OpenStudy (anonymous):

sin 11pi/12 + sin 7pi/12

OpenStudy (anonymous):

Use the trig identity: sin a + sin b = 2 sin [(a + b)/2]* cos [(a - b)/2] sin (a + b)/2 = sin (18Pi/24) = sin (3Pi/4) = V2/2 sin (a - b)/2 - sin (4Pi/24) = sin Pi/6 = V3/2 y = sin 11Pi/12 + sin 7Pi/12 = 2(...)*(...)